OpenStudy (chrisga):

When the function f(x) = 3(5)x is changed to f(x) = 3(5)x + 22, what is the effect?

OpenStudy (chrisga):

There will be no change to the graph because the exponential portion of the function remains the same. The y-intercept is 22 spaces higher. The x-intercept is 22 spaces higher. All input values are moved 22 spaces to the right

OpenStudy (photon336):

Y = m(x)+B M = slope which is change in y over x, and B is your Y intercept. Y = 15x means that the intercept is the origin (0,0) because there's no y intercept it's 0. Y= 15x+22 the graph is shifted up by 22 units on the y axis, because we changed the value of our b = y intercept. To find the value of our y intercept we plug in X = 0 Y= 15(0)+22 (0,22)

OpenStudy (chrisga):

so it B?

OpenStudy (photon336):

That's just the genera equation for a line: Y = m(x)+b I didn't see any answer choices in your question.

OpenStudy (chrisga):

There will be no change to the graph because the exponential portion of the function remains the same. The y-intercept is 22 spaces higher. The x-intercept is 22 spaces higher. All input values are moved 22 spaces to the right

OpenStudy (photon336):

No wait it will be moved up 22 units Y=15X+22 All of your Y values will change

OpenStudy (photon336):

Y intercept is 22 units higher

OpenStudy (chrisga):

ok thank you

OpenStudy (photon336):

@chrisGA absolutely feel free to ask me anything